# ADPO: Anchored Direct Preference Optimization A Unified Framework from Pairwise to Listwise Preferences

### v3

@@ 8 Conclusion
We introduced Anchored Direct Preference Optimization (ADPO), a unified framework that generalizes DPO to soft, listwise settings. We established a key theoretical connection: anchoring imposes an implicit, distribution-space trust region governed by the softmax Fisher metric. This geometric insight provides a principled foundation for understanding preference optimization.
@@
These findings provide clear, actionable guidance: for online RLHF, use dynamic anchors; for offline knowledge transfer, use fixed anchors. By establishing the principle of anchor flexibility, ADPO provides a more robust and versatile tool for aligning models, opening new avenues for adaptive learning strategies in both real-world RL and large-scale language model training. While we have validated our framework on controlled bandits and continuous control, applying this task-aware anchoring strategy to large-language model alignment is a critical and promising direction for future work.
@@ Appendix A Fisher Metric Connection and Relation to KD/TRPO
Fisher metric expansion.
@@
Expanding around $u^{\star}$ such that $\tilde{p}_{\theta}(\cdot;u^{\star})=q$, the second-order Taylor expansion yields
@@
The matrix $\mathrm{Diag}(q)-qq^{\top}$ is exactly the Fisher information of the softmax family, revealing that the anchored KL objective induces a local quadratic form corresponding to the Fisher metric in the distribution space. Therefore, ADPO provides an implicit trust region whose shape is governed by this Fisher geometry.
@@ Appendix B Proofs and Derivations
We prove the three special cases.
(i) Standard DPO (hard labels, no anchoring).
With $q_{ij}\in\{0,1\}$ and $\Delta_{\mathrm{ref}}=0$, the pairwise ADPO loss
$\ell_{ij}=\log\!\left(1+e^{\beta\Delta_{\theta}}\right)-q_{ij}\beta\Delta_{\theta}$
reduces to
$\ell_{ij}=\begin{cases}-\log\sigma(\beta\Delta_{\theta}),&q_{ij}=1,\\[1.99997pt] -\log\sigma(-\beta\Delta_{\theta}),&q_{ij}=0,\end{cases}$
which exactly matches the standard DPO objective [1].
(ii) Non-anchored Bradley–Terry (soft labels, no anchoring).
With $\Delta_{\mathrm{ref}}=0$ and $q_{ij}\in(0,1)$,
$\frac{\partial\ell_{ij}}{\partial\Delta_{\theta}}=\beta\big(\sigma(\beta\Delta_{\theta})-q_{ij}\big).$
Setting the gradient to zero yields $\sigma(\beta\Delta_{\theta})=q_{ij}$, i.e., the Bradley–Terry MLE condition.
(iii) Reward-based soft labels.
If $q_{ij}=\sigma\!\big(\beta_{r}(R_{i}-R_{j})\big)$, then the stationary point satisfies $\sigma\!\big(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}})\big)=\sigma\!\big(\beta_{r}(R_{i}-R_{j})\big)$, so at optimum $\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}})=\beta_{r}(R_{i}-R_{j})$ up to the softmax log-odds, linking policy log-ratios to reward gaps. ∎

### v4

@@ 8 Conclusion
We introduced Anchored Direct Preference Optimization (ADPO), a unified projection framework that reveals fundamental connections across major learning paradigms. By formulating policy learning as an anchored projection—minimizing $\mathrm{KL}(q\|\tilde{p}_{\theta})$ with anchored logits $(s-s^{\mathrm{ref}})/\tau$—we showed that supervised fine-tuning, knowledge distillation, maximum-entropy RL, advantage-weighted methods, and DPO all emerge as special cases through different choices of target distribution, anchor policy, and temperature. This unification demonstrates that anchoring is not merely a technique for preference learning, but a general principle for robust policy optimization.
@@
These findings provide clear, actionable guidance: for online RLHF, use dynamic anchors; for offline knowledge transfer, use fixed anchors; for maximum-entropy RL, use uniform anchors; for advantage-weighted updates, use old-policy anchors. By establishing ADPO as a unifying framework with principled anchor selection, we provide a versatile tool for diverse learning scenarios—from noisy preference optimization to stable knowledge distillation. While we have validated our framework on controlled bandits and continuous control, applying this unified anchoring perspective to large-language model alignment, where different training stages (SFT, RLHF, continual learning) may benefit from different anchor strategies, is a critical and promising direction for future work.
@@ Appendix A Fisher Metric Connection and Relation to KD/TRPO
Fisher metric expansion and Bregman divergence perspective.
@@
Define the convex function $F(u)=A(u)-\langle q,u\rangle$ where $A(u)=\log\sum_{j}e^{u_{j}}$ is the log-partition function. Then the anchored cross-entropy can be written as a Bregman divergence:
$\mathrm{KL}(q\;\|\;\tilde{p}_{\theta})\;=\;F(u)-F(u^{\star})-\langle\nabla F(u^{\star}),u-u^{\star}\rangle,$ (14)
where $u^{\star}$ satisfies $\tilde{p}_{\theta}(\cdot;u^{\star})=q$ (equivalently, $\nabla F(u^{\star})=0$). The Bregman divergence (14) is always non-negative, induces a three-point identity (Pythagorean theorem for Bregman divergences), and provides a unified view of mirror descent and anchored projection updates.
Expanding around $u^{\star}$, the second-order Taylor expansion yields
@@
The matrix $\mathrm{Diag}(q)-qq^{\top}$ is exactly the Fisher information of the softmax family (the Hessian $\nabla^{2}A(u^{\star})$), revealing that the anchored KL objective induces a local quadratic form corresponding to the Fisher metric in the distribution space. Therefore, ADPO provides an implicit trust region whose shape is governed by this Fisher geometry, with connections to mirror descent and Bregman proximal methods.
@@ Appendix B Proofs and Derivations
We prove that ADPO unifies major learning paradigms through different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau$.
(i) Supervised fine-tuning (SFT).
When $\pi_{\mathrm{ref}}$ is uniform (constant logits $s_{i}^{\mathrm{ref}}=c$ for all $i$), the anchored distribution becomes
$\tilde{p}_{\theta}(i\mid S_{x})=\frac{\exp((s_{i}-c)/\tau)}{\sum_{j}\exp((s_{j}-c)/\tau)}=\frac{\exp(s_{i}/\tau)}{\sum_{j}\exp(s_{j}/\tau)}=\mathrm{softmax}_{i}(s_{i}/\tau).$
If $q$ is the data label distribution (one-hot or soft) and $\tau=1$, then $\tilde{p}_{\theta}=\pi_{\theta}$ and $\mathcal{L}_{\mathrm{ADPO}}=\mathbb{E}[\mathrm{CE}(q,\pi_{\theta})]$, the standard cross-entropy loss. For $\tau\neq 1$, the temperature acts as a scaling parameter on the logits (temperature scaling).
(ii) Knowledge distillation (KD).
If $q=\pi_{T}$ (teacher) and $\pi_{\mathrm{ref}}$ is uniform, then with a training temperature $T>1$ we set $\tau=T$ and minimize
$\mathcal{L}_{\mathrm{ADPO}}=\mathbb{E}\!\big[\mathrm{KL}\big(\pi_{T}\;\|\;\mathrm{softmax}(s_{\theta}/T)\big)\big],$
and multiply the loss by $T^{2}$ to recover the classic distillation objective $T^{2}\,\mathrm{CE}\!\big(\pi_{T},\mathrm{softmax}(s_{\theta}/T)\big)$ [22]. Note that $\mathrm{KL}\!\big(\pi_{T}\|\mathrm{softmax}(s_{\theta}/T)\big)=\mathrm{KL}(\pi_{T}\|\pi_{\theta})$ only when $T=1$. If instead $\pi_{\mathrm{ref}}=\pi_{T}$ (anchored KD), we obtain
$\mathcal{L}_{\mathrm{ADPO}}=\mathbb{E}\!\big[\mathrm{KL}\big(\pi_{T}\;\|\;\mathrm{softmax}\!\big((s_{\theta}-s_{T})/\tau\big)\big)\big],$
which performs optimization in teacher-anchored coordinates with an implicit Fisher-metric trust region (see Lemma 3.3).
(iii) Maximum-entropy RL / offline policy projection.
For $q(a\mid s)\propto\exp(Q(s,a)/\alpha)$ and uniform $\pi_{\mathrm{ref}}$, minimizing $\mathcal{L}_{\mathrm{ADPO}}$ yields
$\tilde{p}_{\theta}(a\mid s)=\mathrm{softmax}_{a}(s_{a}/\tau)\propto\exp(Q(s,a)/\alpha),$
so $s_{a}/\tau\approx Q(s,a)/\alpha+\mathrm{const}$. Choosing $\tau=\alpha$ gives the soft-optimal Boltzmann form $\pi\propto\exp(Q/\alpha)$.
For advantage-weighted updates with $q(a\mid s)\propto\exp(A^{\pi_{\mathrm{old}}}(s,a)/\beta)$ and $\pi_{\mathrm{ref}}=\pi_{\mathrm{old}}$, minimizing $\mathcal{L}_{\mathrm{ADPO}}$ yields
$\tilde{p}_{\theta}(a\mid s)=\mathrm{softmax}_{a}\!\left(\frac{s_{a}-s_{a}^{\mathrm{old}}}{\tau}\right)\propto\exp(A^{\pi_{\mathrm{old}}}(s,a)/\beta),$
so $(s_{a}-s_{a}^{\mathrm{old}})/\tau\approx A^{\pi_{\mathrm{old}}}(s,a)/\beta$, giving $\pi_{\theta}(a\mid s)\propto\pi_{\mathrm{old}}(a\mid s)\exp((\tau/\beta)A^{\pi_{\mathrm{old}}}(s,a))$. Choosing $\tau=\beta$ recovers AWR/AWAC/MPO projections.
(iv) DPO (binary comparison form).
For $|S_{x}|=2$ with candidates $\{i,j\}$ and binary $q(i)=1,q(j)=0$ (or Bradley–Terry soft targets $q(i)=\sigma(\beta_{r}\Delta R)$), the ADPO loss becomes
$\mathcal{L}_{\mathrm{ADPO}}=-q(i)\log\tilde{p}_{\theta}(i)-q(j)\log\tilde{p}_{\theta}(j),$
where $\tilde{p}_{\theta}(i)=\sigma((\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau)$ with $\Delta_{\theta}=s_{i}-s_{j}$ and $\Delta_{\mathrm{ref}}=s_{i}^{\mathrm{ref}}-s_{j}^{\mathrm{ref}}$. For uniform $\pi_{\mathrm{ref}}$ (so $\Delta_{\mathrm{ref}}=0$) and hard labels, this reduces to
$\mathcal{L}_{\mathrm{ADPO}}=-\log\sigma(\Delta_{\theta}/\tau),$
which matches the original DPO objective with inverse temperature $1/\tau$. ∎